OpenStudy (sleepyjess):

Let u = <4, 3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

Do you know how to find the magnitude of the vector?

OpenStudy (sleepyjess):

kind of

OpenStudy (anonymous):

Okay, you need to find the magnitude to get the answer, so can you find it or not?

OpenStudy (sleepyjess):

Would I use \(\sf\sqrt{(v_1)^2+ (v_2)^2}\) ?

OpenStudy (anonymous):

Yes

OpenStudy (sleepyjess):

\(\sf\sqrt{4^2~+~3^2}\) \(\sf\sqrt{25}\) 5

OpenStudy (anonymous):

That is right

OpenStudy (anonymous):

That is the right magnitude, so now you just do scalar division